a 5 digit no. abcde is such that the 6 digit no. abcde1 is product of the no. 1abcde and 3.....the sum of the digits of no. abcde is ??? PLS ExPLAIN!!!!

Grade:9

1 Answers

APURV GOEL
13 years ago
Solution :- abcde is a 5 digit no. therefore…… abcde =10000a+1000b+100c+10d+e ……..(1) because, abcde1 = 3*1abcde so, 100000a+10000b+1000c+100d+10e+1 = 3*(100000+10000a+1000b+100c+10d+e) 100000a+10000b+1000c+100d+10e+1 =300000+30000a+3000b+300c+30d+3e 100000a+10000b+1000c+100d+10e-30000a-3000b-300c-30d-3e = 300000-1 70000a+7000b+700c+70d+7e = 299999 7(10000a+1000b+100c+10d+e) = 7(42857) 10000a+1000b+100c+10d+e = 42857 Putting value from 1 we get…………….. abcde = 42857 sum of digits = 4+2+8+5+7 = 26